If a multiple choice test consosts of 6 questions each with 4 possible answers of which only 1 is correct, how many different ways can a student answer the 6 questions
Answer by greenestamps(13203)



4 choices for each of the 6 questions; multiply all the numbers of choices:

4*4*4*4*4*4 = 4^6 = 2^12 = 4096